OpenStudy (anonymous):

Find the antiderivative of (x - 5)/(x^2 - 1)

OpenStudy (anonymous):

partial fractions again :)

OpenStudy (luigi0210):

\[\int\limits \frac{ x-5 }{ (x^2-1) }\]

OpenStudy (luigi0210):

:l

OpenStudy (luigi0210):

Can't you use \[\int\limits \frac{ 1 }{ x^2-a^2 }=\frac{ 1 }{ 2a }\ln |\frac{ x-a }{ x+a }|+C?\]

OpenStudy (anonymous):

yes it is correct, but they likely won't have a formula sheet on the exam. They must be able to reach that without memorizing the formula I believe. This would be by solving for A and B in the case of: \[\int\limits_{}^{} \frac{ x-5 }{ (x-1)(x+1) } dx = \int\limits_{}^{} \frac{ A}{ (x-1)} + \frac{ B }{ x+1 } \]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so i cross multiply next?

OpenStudy (anonymous):

yes you want to make the denominator the same and to solve for A and B using the values on the left

OpenStudy (anonymous):

okay